# Lambda Expected Shortfall

### v1

@@ 1 Introduction
In the landscape of quantitative finance and actuarial science, efficient and robust measurement of risk is paramount. Financial institutions and regulators make use of sophisticated tools to quantify potential losses and manage financial exposures effectively. Among the most widely adopted risk measures are the Value-at-Risk (VaR) and the Expected Shortfall (ES), each with distinct theoretical properties and practical implications. VaR has long served as a standard for risk assessment due to its intuitive interpretability. However, its well-documented limitations, such as the lack of subadditivity and non-convexity for general loss distributions and inability of capturing tail risk (see e.g., Daníelsson et al., 2001; McNeil et al., 2015; Embrechts et al., 2018), have spurred the development of more robust alternatives. ES, also known as the Conditional Value-at-Risk (CVaR), emerged as the most popular alternative, favored for its coherence (Artzner et al., 1999; Acerbi and Tasche, 2002), convexity (Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002), optimization properties (Rockafellar and Uryasev, 2002; Embrechts et al., 2022), and axiomatization via portfolio concentration (Wang and Zitikis, 2021), although it suffers from the lack of elicitability (Gneiting, 2011; Ziegel, 2016; Kou and Peng, 2016; Fissler and Ziegel, 2016).
@@
As a flexible generalization of VaR, the class of Lambda Value-at-Risk ($\Lambda$-VaR) was introduced by Frittelli et al. (2014). The class of $\Lambda$-VaR offers enhanced adaptability for modeling diverse risk preferences and regulatory contexts beyond a fixed confidence level. $\Lambda$-VaR is found to satisfy several useful properties in finance, including monotonicity, cash-subadditivity, elicitability (Bellini and Bignozzi, 2015), robustness (Burzoni et al., 2017), and quasi-star-shapeness (Han et al., 2025). Bellini and Peri (2022) obtained an axiomatic characterization of $\Lambda$-VaR, in particular justifying the choice of $\Lambda$ to be a (weakly) decreasing function. As a risk measure, $\Lambda$-VaR has also been studied from practical aspects such as estimation and backtesting (Hitaj et al., 2018; Corbetta and Peri, 2018), distributionally robust optimizations (Han and Liu, 2025), capital allocations (Ince et al., 2022; Liu, 2025), and optimal insurance problems (Boonen et al., 2025). While $\Lambda$-VaR successfully broadens the scope of VaR, it retains the essential drawbacks of VaR for not being convex and not being able to capture tail risk. A natural remedy for the problem is to introduce an equally flexible generalization of ES as an alternative to $\Lambda$-VaR. However, a suitable way of defining such a risk measure that preserves its desirable properties and strong theoretical foundations has not been found.
@@
This paper addresses this gap by introducing the Lambda Expected Shortfall ($\Lambda$-ES), a novel and theoretically sound risk measure designed to serve as the natural counterpart to $\Lambda$-VaR. One may argue that there are many ways to generalize ES to a class of risk measures parametrized by a function $\Lambda$. A key consideration in ES and its generalization is its consistency with respect to portfolio diversification, modelled via convexity by Föllmer and Schied (2002); Frittelli and Rosazza Gianin (2002). For general risk measures, Cerreia-Vioglio et al. (2011) argued that diversification preferences should be modelled by quasi-convexity, which is equivalent to convexity for monetary risk measures. Keeping this property as our fundamental requirement for a generalization of ES, we find that there is one formulation that has the most advantages, inspired by a recent $\Lambda$-VaR representation result of Han et al. (2025), explained below. For a decreasing function $\Lambda$, we define $\Lambda$-ES of a random variable $X$ by
@@
We demonstrate that, among other potential candidate definitions, $\Lambda$-ES defined as (1) possesses several critical properties (Proposition 2), analogous to those that establish ES as an improved alternative to VaR. Specifically, we prove that, under a mild assumption on $\Lambda$, $\Lambda$-ES defined in (1) is the smallest quasi-convex and law-invariant risk measure that dominates $\Lambda$-VaR (Theorem 2). This generalization is a significant theoretical advancement, extending the dominance relationship from VaR versus ES (Delbaen, 2012; Föllmer and Schied, 2016) to the more flexible $\Lambda$-VaR versus $\Lambda$-ES framework. As a by-product, we obtain a new result on the domination of ES over VaR (Theorem 1) that is stronger than several classic results in the literature.
Beyond the foundational definition and properties of $\Lambda$-ES, which are the topics of Section 3, we proceed to conduct a comprehensive analysis of this new class of risk measures. In Section 4, we obtain a dual representation of $\Lambda$-ES (Theorem 3), offering deeper insights into its theoretical structure and connections to quasi-convex cash-subadditive risk measures. In Section 5, we explore the properties of $\Lambda$-ES in optimization problems, both as an objective function to minimize and as a constraint to impose, and analyze various forms of convexity in relevant reformulations of ES optimization problems. As standard in the risk measures literature, the main results are formulated on the space $L^{\infty}$ of essentially bounded random variables. In Section 6, results are naturally extended to the space $L^{1}$ of integrable random variables, sometimes under slightly stronger assumptions. Section 7 concludes the paper. Some alternative potential formulations for $\Lambda$-ES are discussed in Appendix A, demonstrating why our proposed definition is the most robust, theoretically consistent, and desirable for risk management applications.
@@ 2.1 Risk measures
Let $L^{0}$ be the space of all random variables on an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, $L^{1}$ be the space of all random variables with finite mean, and $L^{\infty}$ be the set of all essentially bounded random variables. Write $\overline{\mathbb{R}}=[-\infty,\infty]$ and $\mathbb{R}_{+}=[0,\infty)$. For any $n\in\mathbb{N}$, denote by $n]=\{1,\dots,n\}$. For any $x,y\in\overline{\mathbb{R}}$, write $x\wedge y=\min\{x,y\}$, $\vee y=\max\{x,y\}, x+=x$$_{+}=x\vee 0$, and $x_{-}=x\wedge 0$. For any function $f:\mathbb{R}\to\mathbb{R}$ and $x\in\mathbb{R}$, we write $f(x-)=\lim_{y\uparrow x}f(y)$ and $f(x+)=\lim_{y\downarrow x}f(y)$, if they exist.
We start with risk measures that are used to quantify risks. A risk measure is a mapping $\rho:\mathcal{X}\to\overline{\mathbb{R}}$, where $\mathcal{X}$ is the space of random variables where $\rho$ is defined. Below, we list several properties that a risk measure may satisfy. The functional $\rho$ is called a monetary risk measure if it satisfies
@@
As standard considerations for risk measures (e.g., Artzner et al., 1999), monotonicity naturally means that increasing the amount of loss will lead to more risk of the financial position, whereas cash additivity ensures that $\rho(X)$ is the total capital required to be added to the financial position to make it acceptable given the loss of $X\in\mathcal{X}$. It is also natural to assume a monetary risk measure satisfies
@@
A monetary risk measure is called coherent (Artzner et al., 1999) if it further satisfies [^1] Whenever convexity or subadditivity is discussed, the range of $\rho$ includes at most one of $\infty$ and -$\infty$ to avoid $\infty$-$\infty$.
@@
whereas a monetary risk measure is called a convex risk measure (Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002) if it further satisfies
@@
Convexity is motivated by diversification effects in risk measurement. We refer to Föllmer and Schied (2016) for a comprehensive review of properties of monetary risk measures. Beyond the monetary framework, El Karoui and Ravanelli (2009) argued that cash additivity fails to incorporate the ambiguity of interest rates and thus proposed the property of
@@
Many commonly used convex risk measures (such as the Expected Shortfall defined below) also satisfy law invariance and concavity with respect to distribution mixtures, defined below. Let $\mathcal{M}_{c}(\mathbb{R})$ denote the set of compactly supported distributions on $\mathbb{R}$.
  • Law-invariance: $\rho(X)=\rho(Y)$ for all $X,Y\in\mathcal{X}$ with the same distribution.
  • Concavity (resp. quasi-concavity) in mixtures: $\rho$ is law-invariant and the function $F\mapsto\rho(X_{F})$ on $\mathcal{M}_{c}(\mathbb{R})$ is concave (resp. quasi-concave), where $X_{F}$ is a random variable with distribution $F\in\mathcal{M}_{c}(\mathbb{R})$.
@@
  • SSD-consistency: $\rho(X)\geqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ and $X\succeq_{\rm icx}Y$. [^2] Here, SSD represents second-order stochastic dominance. For $X,Y\in\mathcal{X}$, we say that $X$ dominates $Y$ in increasing convex order, denoted by $X\succeq_{\rm icx}Y$, if $\mathbb{E}[f(X)]\geqslant\mathbb{E}[f(Y)]$ for all increasing and convex functions $f:\mathbb{R}\to\mathbb{R}$.
@@
For the financial interpretation and characterization of the properties introduced above, we refer to Föllmer and Schied (2002, convexity), Cerreia-Vioglio et al. (2011, quasi-convexity), Mao and Wang (2020, SSD-consistency), and Wang et al. (2020, concavity in mixtures).
@@ 2.2 VaR and Lambda VaR
The Value-at-Risk (VaR) at level $\alpha\in[0,1]$ is defined as the left-quantile, namely,
@@
Note that for all $X\in L^{0}$, we have $\mathrm{VaR}_{0}(X)=$-$\infty$ and $\mathrm{VaR}_{1}(X)$ is the essential supremum of $X$. Similarly, the upper Value-at-Risk ($\mathrm{VaR}^{+}$) at level $\alpha\in[0,1]$ is defined as the right-quantile:
@@
Using these formulations, $\mathrm{VaR}^{+}_{0}(X)$ is the essential infimum of $X$ and $\mathrm{VaR}^{+}_{1}(X)=\infty$. Moreover, $\mathrm{VaR}_{\alpha}(X),\mathrm{VaR}^{+}_{\alpha}(X)\in\mathbb{R}$ for any $\alpha\in(0,1)$. Both versions of VaR satisfy monotonicity, cash additivity, positive homogeneity, law invariance, and quasi-concavity in mixtures, but not quasi-convexity, concavity in mixtures, or SSD-consistency. Indeed, the latter three properties are equivalent for the class of distortion risk measures; see Wang et al. (2020, Theorem 3).
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. Throughout, all terms like “increasing” and “decreasing” are in the weak sense. The $\Lambda$-Value-at-Risk ($\Lambda$-VaR, or $\Lambda$-quantile), denoted by $\mathrm{VaR}_{\Lambda}:L^{0}\to\overline{\mathbb{R}}$, is defined as
@@
Frittelli et al. (2014) originally introduced $\Lambda$-VaR focusing on the case that $\Lambda$ is increasing. Bellini and Bignozzi (2015) showed that $\Lambda$-VaR with a decreasing $\Lambda$ satisfies elicitability, and it is not true for increasing $\Lambda$ (Burzoni et al., 2017). A more decisive result is the axiomatic justification of Bellini and Peri (2022) for using a decreasing function $\Lambda$. Han et al. (2025) further showed that $\Lambda$-VaR with a decreasing $\Lambda$ is cash-subadditive and hence $L^{\infty}$-continuous, but with an increasing $\Lambda$ even $L^{\infty}$-continuity fails. For these reasons, we focus on the case of decreasing $\Lambda$.
The quantile has two versions, and so does $\Lambda$-VaR. We define the upper $\Lambda$-Value-at-Risk ($\Lambda$-$\mathrm{VaR}^{+}$ or upper $\Lambda$-quantile), denoted by $\mathrm{VaR}^{+}_{\Lambda}:L^{0}\to\overline{\mathbb{R}}$, as
@@
If $\Lambda$ is a constant $\alpha\in[0,1]$ then $\mathrm{VaR}_{\Lambda}=\mathrm{VaR}_{\alpha}$ (resp. $\mathrm{VaR}^{+}_{\Lambda}=\mathrm{VaR}^{+}_{\alpha}$), which is the left- (resp. right-) quantile at level $\alpha$. The $\Lambda$-VaR is monotone, but not cash additive or positively homogeneous, thus losing some usual properties of VaR.
@@ 2.3 Expected Shortfall
We present the standard risk measure in banking regulation, the Expected Shortfall (ES), via several different formulations. First, as the most standard definition, ES at level $\alpha\in[0,1]$ is formally defined as the mapping $\mathrm{ES}_{\alpha}:L^{\infty}\to\mathbb{R}$ given by
@@
and $\mathrm{ES}_{1}(X)=\mathrm{VaR}_{1}(X)$. Note that $\mathrm{ES}_{0}=\mathbb{E}$ and the definition of $\mathrm{ES}_{\alpha}$ in (2) can be easily extended to $L^{1}$, and we will discuss risk measures on $L^{1}$ in Section 6. An ES satisfies all properties listed in Section 2.1.
Second, as shown by Rockafellar and Uryasev (2002), for $\alpha\in[0,1]$, $\mathrm{ES}_{\alpha}$ can be equivalently formulated by
@@
where we use the convention $0/0=0$ and $x/0=\infty$ for $x>0$. The representation (3) connects to VaR via
@@
We may call (3) and (6) the Rockafellar–Uryasev (RU) relation for VaR and ES.
Third, it is known that the risk measure $\mathrm{ES}_{\alpha}$ is the smallest law-invariant coherent risk measure dominating $\mathrm{VaR}_{\alpha}$ (Delbaen, 2012, Theorem 52). Convexity is important and relevant for risk management, and for this reason, ES is regarded as an improvement of VaR. In the next result, we show that $\mathrm{ES}_{\alpha}$ is also the smallest mapping dominating $\mathrm{VaR}_{\alpha}$ satisfying quasi-convexity and law invariance. As far as we know, this result is new, and it is based on a VaR-ES asymptotic equivalence result of Wang and Wang (2015) and a result in Embrechts et al. (2015) on the sum of negatively dependent sequences. Throughout the paper, we write $\rho\geqslant\widetilde{\rho}$ for mappings $\rho:\mathcal{X}\to\overline{\mathbb{R}}$ and $\widetilde{\rho}:\widetilde{\mathcal{X}}\to\overline{\mathbb{R}}$ to represent the dominance of $\rho$ over $\widetilde{\rho}$ on their common domain (i.e., $\rho(X)\geqslant\widetilde{\rho}(X)$ for all $X\in\mathcal{X}\cap\widetilde{\mathcal{X}}$), and typically we have either $X\subseteq\widetilde{\mathcal{X}}$ or $\widetilde{X}\subseteq\mathcal{X}$.
@@
For any $\alpha\in(0,1]$,
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (7)
For any $\alpha\in[0,1)$,
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (8)
@@
(i) We first prove (7). Let $\rho$ be quasi-convex and law-invariant satisfying $\rho\geqslant\mathrm{VaR}_{\alpha}$. If $\alpha=1$, it is clear that (7) holds, because $\mathrm{VaR}_{1}$ is quasi-convex and law-invariant. Next, suppose $\alpha\in(0,1)$. For any $X\in L^{\infty}$ with distribution $F$, we have
@@
For $\alpha=0$ and $X\in L^{\infty}$ with distribution $F$, write $M=\mathrm{VaR}_{1}(X)-\mathrm{VaR}^{+}_{0}(X)$. For any $n\in\mathbb{N}$, by Corollary A.3 of Embrechts et al. (2015), there exist $\widetilde{X}_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F$, $i\in[n]$, such that
$\left|\frac{1}{n}\sum^{n}_{i=1}\widetilde{X}_{i}-\mathbb{E}[X]\right|\leqslant\frac{M}{n}.$
Hence,
$\frac{1}{n}\sum^{n}_{i=1}\widetilde{X}_{i}\geqslant\mathbb{E}[X]-\frac{M}{n}.$
It yields that

### v2

@@ 1 Introduction
Financial institutions and regulators make use of sophisticated tools to quantify potential losses and manage financial exposures effectively. Among the most widely adopted risk measures are the Value-at-Risk (VaR) and the Expected Shortfall (ES), each with distinct theoretical properties and practical implications. VaR has long served as a standard for risk assessment due to its intuitive interpretability. However, its well-documented limitations, such as the lack of subadditivity and non-convexity for general loss distributions and inability of capturing tail risk (see e.g., Daníelsson et al., 2001; McNeil et al., 2015; Embrechts et al., 2018), have spurred the development of more robust alternatives. ES, also known as the Conditional Value-at-Risk (CVaR), emerged as the most popular alternative, with desirable features such as coherence (Artzner et al., 1999; Acerbi and Tasche, 2002), convexity (Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002), optimization properties (Rockafellar and Uryasev, 2002; Embrechts et al., 2022), and axiomatization via portfolio concentration (Wang and Zitikis, 2021), although it suffers from the lack of elicitability (Gneiting, 2011; Ziegel, 2016; Kou and Peng, 2016; Fissler and Ziegel, 2016).
@@
As a flexible generalization of VaR, the class of Lambda Value-at-Risk ($\Lambda$-VaR) was introduced by Frittelli et al. (2014). The class of $\Lambda$-VaR offers enhanced adaptability for modeling diverse risk preferences and regulatory contexts beyond a fixed confidence level. $\Lambda$-VaR is found to satisfy several useful properties in finance, including monotonicity, cash subadditivity, elicitability (Bellini and Bignozzi, 2015), robustness (Burzoni et al., 2017), and quasi-star-shapeness (Han et al., 2025). Bellini and Peri (2022) obtained an axiomatic characterization of $\Lambda$-VaR, in particular justifying the choice of $\Lambda$ to be a (weakly) decreasing function. As a risk measure, $\Lambda$-VaR has also been studied from practical aspects such as estimation and backtesting (Hitaj et al., 2018; Corbetta and Peri, 2018), distributionally robust optimizations (Han and Liu, 2025), capital allocations (Ince et al., 2022; Liu, 2025), and optimal insurance problems (Boonen et al., 2025). While $\Lambda$-VaR successfully broadens the scope of VaR, it retains the essential drawbacks of VaR for not being convex and not being able to capture tail risk. A natural remedy for the problem is to introduce an equally flexible generalization of ES as an alternative to $\Lambda$-VaR. A suitable way of defining such a risk measure that preserves its desirable properties and strong theoretical foundations has not been found.
@@
This paper addresses this gap by introducing the Lambda Expected Shortfall ($\Lambda$-ES), a natural counterpart to $\Lambda$-VaR. There are many potential ways to generalize ES to a class of risk measures parametrized by a function $\Lambda$. A key consideration in ES and its generalization is its consistency with respect to portfolio diversification, modelled via convexity by Föllmer and Schied (2002); Frittelli and Rosazza Gianin (2002). For general risk measures, Cerreia-Vioglio et al. (2011) argued that diversification preferences should be modelled by quasi-convexity, which is equivalent to convexity for monetary risk measures. Keeping this property as our fundamental requirement for a generalization of ES, we find that there is one formulation that has the most advantages, inspired by a recent $\Lambda$-VaR representation result of Han et al. (2025). For a decreasing function $\Lambda$, we define $\Lambda$-ES of a random variable $X$ by
@@
We demonstrate that $\Lambda$-ES defined in (1) possesses several critical properties (Proposition 2), analogous to those that establish ES as an improved alternative to VaR. More importantly, based on a new result on the domination of ES over VaR (Theorem 1), we show that $\Lambda$-ES is the smallest quasi-convex and law-invariant risk measure that dominates $\Lambda$-VaR (Theorem 2). This result extends the classic dominance between VaR and ES (Delbaen, 2012; Föllmer and Schied, 2016).
Beyond the foundational definition and properties of $\Lambda$-ES, which are the topics of Section 3, we proceed to conduct a comprehensive analysis of this new class of risk measures. In Section 4, we obtain a dual representation of $\Lambda$-ES (Theorem 3), offering deeper insights into its theoretical structure and connections to quasi-convex cash-subadditive risk measures. In Section 5, we explore the properties of $\Lambda$-ES in optimization problems, both as an objective function to minimize and as a constraint to impose, and analyze various forms of convexity in relevant reformulations of ES optimization problems. In Section 6, results are naturally extended to the space $L^{1}$ of integrable random variables, sometimes under slightly stronger assumptions. Section 7 concludes the paper. Some alternative potential formulations for $\Lambda$-ES are discussed in Appendix A, demonstrating why our proposed definition is the most robust, theoretically consistent, and desirable for risk management applications.
@@ 2.1 Risk measures
Let $L^{0}$ be the space of all random variables on an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, $L^{1}$ be the space of all random variables with finite mean, and $L^{\infty}$ be the set of all essentially bounded random variables. Write $\overline{\mathbb{R}}=[-\infty,\infty]$ and $\mathbb{R}_{+}=[0,\infty)$. For any $n\in\mathbb{N}$, denote by $n]=\{1,\dots,n\}$. For any $x,y\in\overline{\mathbb{R}}$, write $x\wedge y=\min\{x,y\}$, $\vee y=\max\{x,y\}, x+=x$$_{+}=x\vee 0$, and $x_{-}=x\wedge 0$. For any function $f:\mathbb{R}\to\mathbb{R}$ and $x\in\mathbb{R}$, we write $f(x-)=\lim_{y\uparrow x}f(y)$ and $f(x+)=\lim_{y\downarrow x}f(y)$, if they exist. Let $\mathcal{M}_{c}(\mathbb{R})$ denote the set of compactly supported distributions on $\mathbb{R}$.
We start with risk measures that are used to quantify risks. A risk measure is a mapping $\rho:\mathcal{X}\to\overline{\mathbb{R}}$, where $\mathcal{X}$ is the space of random variables where $\rho$ is defined. Below, we list several common properties that a risk measure may satisfy, and their financial interpretation is well documented in the literature (e.g., Artzner et al., 1999; Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002). The risk measure $\rho$ is called a monetary risk measure if it satisfies
@@
A monetary risk measure is often required to satisfy
@@
A monetary risk measure is called coherent if it further satisfies [^1] Whenever convexity or subadditivity is discussed, the range of $\rho$ includes at most one of $\infty$ and -$\infty$ to avoid $\infty$-$\infty$.
@@
whereas a monetary risk measure is called a convex risk measure if it further satisfies
@@
Convexity is motivated by diversification effects in risk measurement. To incorporate non-constant interest rates, El Karoui and Ravanelli (2009) relaxes cash additivity to
@@
Many commonly used convex risk measures (such as the Expected Shortfall defined below) also satisfy law invariance and concavity with respect to distribution mixtures.
  • Law invariance: $\rho(X)=\rho(Y)$ for all $X,Y\in\mathcal{X}$ with the same distribution.
  • Concavity (resp. quasi-concavity) in mixtures: $\rho$ is law invariant and the function $F\mapsto\rho(X_{F})$ on $\mathcal{M}_{c}(\mathbb{R})$ is concave (resp. quasi-concave), where $X_{F}$ is a random variable with distribution $F\in\mathcal{M}_{c}(\mathbb{R})$.
@@
  • SSD-consistency: $\rho(X)\geqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ and $X\succeq_{\rm icx}Y$. [^2] Here, SSD represents second-order stochastic dominance. For $X,Y\in\mathcal{X}$, we say that $X$ dominates $Y$ in increasing convex order, denoted by $X\succeq_{\rm icx}Y$, if $\mathbb{E}[f(X)]\geqslant\mathbb{E}[f(Y)]$ for all increasing and convex functions $f:\mathbb{R}\to\mathbb{R}$. SSD-consistent monetary risk measures are characterized by Mao and Wang (2020).
@@ 2.2 VaR and Lambda VaR
The Value-at-Risk (VaR) at level $\alpha\in[0,1]$ is defined as the left quantile, namely,
@@
Similarly, the upper Value-at-Risk ($\mathrm{VaR}^{+}$) at level $\alpha\in[0,1]$ is defined as the right quantile:
@@
Using these formulations, for any $X\in L^{0}$, $\mathrm{VaR}_{0}(X)=$-$\infty$, $\mathrm{VaR}^{+}_{1}(X)=\infty$, $\mathrm{VaR}_{1}(X)$ is the essential supremum of $X$, and $\mathrm{VaR}^{+}_{0}(X)$ is the essential infimum of $X$. Moreover, $\mathrm{VaR}_{\alpha}(X),\mathrm{VaR}^{+}_{\alpha}(X)\in\mathbb{R}$ for any $\alpha\in(0,1)$. Both versions of VaR satisfy monotonicity, cash additivity, positive homogeneity, law invariance, and quasi-concavity in mixtures, but not quasi-convexity, concavity in mixtures, or SSD-consistency.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. Throughout, all terms of “increasing” and “decreasing” are in the weak sense. The $\Lambda$-Value-at-Risk ($\Lambda$-VaR, or $\Lambda$-quantile), denoted by $\mathrm{VaR}_{\Lambda}:L^{0}\to\overline{\mathbb{R}}$, is defined as
@@
Frittelli et al. (2014) originally introduced $\Lambda$-VaR focusing on the case that $\Lambda$ is increasing. Bellini and Bignozzi (2015) showed that $\Lambda$-VaR with a decreasing $\Lambda$ satisfies elicitability, and it is not true for increasing $\Lambda$ (Burzoni et al., 2017). A more decisive result is the axiomatic justification of Bellini and Peri (2022) for using a decreasing function $\Lambda$. Han et al. (2025) further showed that $\Lambda$-VaR with a decreasing $\Lambda$ is cash subadditive and hence $L^{\infty}$-continuous, but with an increasing $\Lambda$ even $L^{\infty}$-continuity fails. For these reasons, our study focuses on the case of decreasing $\Lambda$.
VaR has two versions, and so does $\Lambda$-VaR. We define the upper $\Lambda$-VaR, denoted by $\mathrm{VaR}^{+}_{\Lambda}:L^{0}\to\overline{\mathbb{R}}$, as
@@
If $\Lambda$ is a constant $\alpha\in[0,1]$ (written $\Lambda\equiv\alpha$), then $\mathrm{VaR}_{\Lambda}=\mathrm{VaR}_{\alpha}$ and $\mathrm{VaR}^{+}_{\Lambda}=\mathrm{VaR}^{+}_{\alpha}$. The risk measure $\Lambda$-VaR is monotone, but not cash additive or positively homogeneous, thus losing some usual properties of VaR.
Han et al. (2025, Theorem 1) gives a representation of $\Lambda$-VaR, which will be useful for our study. Below, we state the result and extend it to $\Lambda$-$\mathrm{VaR}^{+}$.
Proposition 1.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. The risk measures $\mathrm{VaR}_{\Lambda}$ and $\mathrm{VaR}^{+}_{\Lambda}$ admit the following representations:
Proof.
Equation (2) holds directly by Theorem 1 of Han et al. (2025) for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$ that are not constantly 0. For $\Lambda\equiv 0$, we have
Thus (2) holds for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$. To see (3), using some standard relations between quantiles and distribution functions, we have
$=\sup\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)\geqslant x\}=\sup_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x\right\},$
and similarly,
$\mathrm{VaR}^{+}_{\Lambda}(X)=\inf\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)<x\}=\inf_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x\right\}.$
The proof is complete. ∎
@@ 2.3 Expected Shortfall
The standard risk measure in banking regulation, the Expected Shortfall (ES), can be defined via a few different formulations. First, as the most standard definition, ES at level $\alpha\in[0,1]$ is defined as the mapping $\mathrm{ES}_{\alpha}:L^{\infty}\to\mathbb{R}$ given by
@@
and $\mathrm{ES}_{1}(X)=\mathrm{VaR}_{1}(X)$. Note that $\mathrm{ES}_{0}=\mathbb{E}$ and the definition of $\mathrm{ES}_{\alpha}$ in (4) can be easily extended to $L^{1}$. We will discuss risk measures on $L^{1}$ in Section 6. An ES satisfies all properties listed in Section 2.1. Second, as shown by Rockafellar and Uryasev (2002), for $\alpha\in[0,1]$, $\mathrm{ES}_{\alpha}$ can be equivalently formulated by
@@
where we set $0/0=0$ and $x/0=\infty$ for $x>0$. The representation (5) connects to VaR via
@@
We call (5) and (8) the Rockafellar–Uryasev (RU) formulas for VaR and ES. Third, it is known that the risk measure $\mathrm{ES}_{\alpha}$ is the smallest law-invariant coherent risk measure dominating $\mathrm{VaR}_{\alpha}$ (Delbaen, 2012, Theorem 52). Convexity is important and relevant for risk management, and for this reason, ES is regarded as an improvement of VaR. In the next result, we show that $\mathrm{ES}_{\alpha}$ is also the smallest mapping dominating $\mathrm{VaR}_{\alpha}$ satisfying quasi-convexity and law invariance. As far as we know, this result is new, and it is based on a VaR-ES asymptotic equivalence result of Wang and Wang (2015) and a result in Embrechts et al. (2015) on the sum of negatively dependent sequences. Throughout the paper, we write $\rho\geqslant\widetilde{\rho}$ for mappings $\rho:\mathcal{X}\to\overline{\mathbb{R}}$ and $\widetilde{\rho}:\widetilde{\mathcal{X}}\to\overline{\mathbb{R}}$ to represent the dominance of $\rho$ over $\widetilde{\rho}$ on their common domain (i.e., $\rho(X)\geqslant\widetilde{\rho}(X)$ for all $X\in\mathcal{X}\cap\widetilde{\mathcal{X}}$), and typically we have either $\mathcal{X}\subseteq\widetilde{\mathcal{X}}$ or $\widetilde{\mathcal{X}}\subseteq\mathcal{X}$.
@@
The following equalities hold:
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law invariant}\},~\alpha\in(0,1].$ (9)
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law invariant}\},~\alpha\in[0,1).$ (10)
@@
(i) We first prove (9). Let $\rho$ be quasi-convex and law invariant satisfying $\rho\geqslant\mathrm{VaR}_{\alpha}$. If $\alpha=1$, it is clear that (9) holds, because $\mathrm{VaR}_{1}$ is quasi-convex and law invariant. Next, suppose $\alpha\in(0,1)$. For any $X\in L^{\infty}$ with distribution $F$, we have
@@
For $\alpha=0$ and $X\in L^{\infty}$ with distribution $F$, write $M=\mathrm{VaR}_{1}(X)-\mathrm{VaR}^{+}_{0}(X)$. For any $n\in\mathbb{N}$, by Corollary A.3 of Embrechts et al. (2015), there exist $\widetilde{X}_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F$, $i\in[n]$, such that $|\sum^{n}_{i=1}\widetilde{X}_{i}/n-\mathbb{E}[X]|\leqslant M/n.$ Hence, $\sum^{n}_{i=1}\widetilde{X}_{i}/n\geqslant\mathbb{E}[X]-M/n.$ It yields that